OpenStudy (anonymous):

Which equation implies that A and B are independent events? P(A B) = P(A) P(B) P(A) = P(B) P(B|A) = P(A) P(A B) = P(A) P (B) P(A|B) = P(B|A)

OpenStudy (ybarrap):

It's hard to read your options, but P(A|B) = P(A) implies A is independent

OpenStudy (anonymous):

sorry, but the first one is P(A) over P(B)

OpenStudy (ybarrap):

Over B? Like this P(A)/P(B) or P(A\B) or P(A|B) ?

OpenStudy (anonymous):

for independent events \[p(a \cap b )= p(a)*p(b)\]

OpenStudy (ybarrap):

You can also have independent events if P(AB)=P(A)P(B) or P(B|A)=P(B)

OpenStudy (anonymous):

The intersection symbol is missing from them

OpenStudy (ybarrap):

There is a Section called "Latex Practicing" that will help you to quickly come up to speed with writing symbols that is very helpful. But I think you have the info you need. Use AND instead of \(\cap\) and use OR instead of \(\cup\) for now .